Tuesday 19th Oct
A couple of problems to sort out at CRC today - so a trip with P & O with shovel, pick and tools in buckets.
The Signal wires opposite the signal box where they disappear into the tube to pass under the foot crossing have become snagged in weeds and ballast so we have cut back and cleared the problem.
The Signal wire for signal 1 has become too tight and is at the limit of its adjuster. Malcolm has made a link to extend it in the locking room but first we must tie off the wire outside to give some slack to allow disconnection and insert the link.
A bit awkward getting in between the wires, pulleys and levers but the link has been inserted and the adjuster now has about 6" to play with and the signal showing a good "off"
While we were in the locking room we gave all the pulleys and angle crank pivots an oiling.
A picture from Peter W concerning the refurbishment of a token instrument by Andy D. There are a lot of damaged tokens which have been cut up to be sure that they cannot be used. Hope to get some pictures and a bit more detail of this instrument next week